Given the | x - 3 | = 5 equation, what is the average of its solutions?
Answer by ankor@dixie-net.com(22740)

Given the | x - 3 | = 5 equation, what is the average of its solutions?
:
|x-3| = 5
Remove the absolutes, two solutions
x - 3 = 5
x = 5 + 3
x = 8
and
x - 3 = -5
x = -5 + 3
x = -2
:
Find the difference between -2 and 8
dif = 8 - (-2)
dif = 8 + 2
dif = 10
Subtract half of 10 from 8
av = +3